OpenStudy (anonymous):

If sinA + sinB=1 and sinA-sinB=1, find A.

OpenStudy (anonymous):

sinA + sinB=1 and sinA-sinB=1 Then, adding both the equation gives: 2sinA =2 sinA=1

OpenStudy (anonymous):

Now,

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

i thought it will be like sum and difference. you squared the first term - the squared of the 2nd term

OpenStudy (anonymous):

No..... I just added both the equations

OpenStudy (anonymous):

it just depend in the equation :)) okay thanks.. i got it :) can i know your secret? haha

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

And a-b = a+ b = 1 has to be b = 0

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

sauravshakya just showed you why, but it should be clear from just looking at it.

OpenStudy (anonymous):

anyway. thanks to your help :) i'll just ask a question again later :)

OpenStudy (anonymous):

ur welcome